# Some Bayesian Perspectives on Clinical Trials

### v1

@@ 4.2 Adaptive enrichment
When a treatment benefits only a biomarker-defined subgroup, an enrichment design modifies enrollment criteria based on interim data. Starting with broad enrollment, the trial evaluates treatment effects within pre-specified subgroups. If evidence accumulates that the treatment is ineffective in certain subgroups, enrollment is restricted to the responsive population.
The decision at each interim analysis—whether to continue broad enrollment, restrict to a subgroup, or stop entirely—is a multi-armed bandit problem where the “arms” are enrollment strategies rather than treatments. The posterior predictive probability of eventual success in each subgroup guides the decision.
@@ 5.1 ECMO: the role of priors with minimal data
The 1985 trial of extracorporeal membrane oxygenation (ECMO) for newborns with persistent pulmonary hypertension [Bartlett et al., 1985] used a randomized play-the-winner design [Wei and Durham, 1978]: adaptive randomization that favors the more successful treatment. The first patient was randomized to ECMO and survived; the second to conventional medical therapy (CMT) and died; the remaining ten were all randomized to ECMO and all survived. Final tally: 11/11 ECMO survivors, 0/1 CMT survivors.
Ware [1989] discussed the statistical challenges; Kass and Greenhouse [1989] provided the Bayesian perspective. With a uniform prior $\text{Beta}(1,1)$ on each treatment’s survival probability, the posteriors are:
@@
The posterior probability $P(\theta_{\text{ECMO}}>\theta_{\text{CMT}}\mid\text{data})$ can be computed exactly via Thompson’s formula (3) and exceeds 0.99.
The critical issue is sensitivity to the prior. With only one control patient, the posterior for $\theta_{\text{CMT}}$ is dominated by the prior. Historical data suggested approximately 20% survival with CMT from large observational series. Encoding this as $\text{Beta}(4,16)$—equivalent to 20 prior patients with 4 survivors—yields $\theta_{\text{CMT}}\mid\text{data}\sim\text{Beta}(4,17)$ with posterior mean 0.19, and $P(\theta_{\text{ECMO}}>\theta_{\text{CMT}}\mid\text{data})$ remains above 0.99. The conclusion is robust because the ECMO evidence (11/11) is overwhelming; the prior matters only for the magnitude of the estimated difference, not the direction.
Berry [1989] emphasized that the Bayesian framework naturally handles the controversial stopping rule: under the likelihood principle, inference depends on the data actually observed, not on the stopping rule that generated them. A p-value for the same data depends on whether the investigator planned to stop after 12 patients, planned to stop after one CMT death, or had no fixed plan—the Bayesian posterior does not.
The ECMO trial was reanalyzed in the modern EOLIA trial for adult ARDS [Goligher et al., 2018]. The frequentist p-value was 0.09 (not significant at the conventional 0.05 level), but Bayesian analyses with various priors showed 88–99% posterior probability that ECMO reduces mortality. This illustrates how the Bayesian framework provides more nuanced and actionable conclusions than a binary significant/not-significant determination.
@@ 5.2 CALGB 49907: predictive probability monitoring
The Cancer and Leukemia Group B trial 49907 tested whether capecitabine could replace standard chemotherapy in women with early-stage breast cancer aged 65 or older [Muss et al., 2009]. The planned sample size ranged from 600 to 1800 patients, with Bayesian predictive probability monitoring for futility and noninferiority at pre-specified interim analyses.
Capecitabine would be declared noninferior if the hazard ratio for disease recurrence or death exceeded 0.8046 (corresponding to a 5-year relapse-free survival of 53% vs. 60% for standard therapy). At each interim analysis, the predictive probability was computed: given current data, what is the probability that future follow-up will yield a conclusive result?
At the first interim analysis after 600 patients, the hazard ratio was 0.53 (capecitabine markedly inferior). The Bayesian probability that this ratio was less than 0.8046 was 96%, exceeding the futility threshold of 80%. Enrollment was discontinued. After an additional year of follow-up, the hazard ratio was 2.09 (95% CI: 1.38–3.17; $P<0.001$), confirming that standard chemotherapy is superior.
The trial achieved its scientific goal with 633 patients instead of the 1800 maximum—a savings of 65% in sample size and years of accrual time. This was the first fully Bayesian randomized trial published in the New England Journal of Medicine [Muss et al., 2009]. The design used noninformative priors for treatment effects; the Bayesian contribution was entirely in the decision framework (predictive probabilities for stopping), not in the priors.
@@ 5.3 Platform trials and the Bayesian Time Machine
Saville et al. [2022] proposed the Bayesian Time Machine: a hierarchical model that smooths control-arm response rates across time, allowing analyses to use all control data while adjusting for temporal drift. The model includes time effects as random effects with a smoothing prior that penalizes abrupt changes, so that closer time intervals are modeled with similar response rates.
In simulations, the Bayesian Time Machine provides estimates with smaller mean squared error than either concurrent-only or pooled-controls analyses. The smoothing prior acts as a compromise: when temporal drift is absent, it borrows efficiently across all time periods; when drift is present, it limits borrowing to adjacent periods. The I-SPY 2 platform trial for neoadjuvant breast cancer employed this approach across 23 experimental therapies, of which nine graduated to Phase III based on exceeding an 85% Bayesian predictive probability of confirmatory trial success [Berry, 2025].
@@ 6 The Regulatory Landscape
The FDA’s January 2026 draft guidance on Bayesian methodology [Food and Drug Administration, 2026] represents a significant shift from philosophical debate to practical specification. The guidance requires:
  • Pre-specification of Bayesian success criteria and decision thresholds.
  • Evaluation of operating characteristics through simulation, including type I error under the null hypothesis.
  • Justification of prior distributions, with quantification of their influence (e.g., effective sample size).
  • Sensitivity analyses examining robustness to prior specification and prior-data conflict.
@@
The guidance identifies specific settings where Bayesian methods have supported successful submissions: borrowing from previous trials (e.g., REBYOTA for recurrent C. difficile), augmenting concurrent controls with external data (e.g., GBM AGILE, Precision Promise), pediatric extrapolation (e.g., empagliflozin, linagliptin for pediatric type 2 diabetes), and dose-finding trials in oncology.
A common misconception, noted even in the draft ICH E20 guideline on adaptive trials, equates Bayesian methods primarily with informative priors for borrowing. The majority of successful Bayesian adaptive trials use minimally informative priors overwhelmed by accumulating data. The benefits come from the coherent inferential framework, the interpretability of posterior probability statements, and predictive probabilities—which have no frequentist analog.
@@ 7 Discussion
The three themes of this paper—prior specification, sequential design, and decision-theoretic optimization—are not independent ideas that happen to share a Bayesian label. They are consequences of a single principle: probability is the language of uncertainty, and decisions under uncertainty should maximize expected utility.
When this principle is applied to clinical trials, priors are not optional nuisance parameters but expressions of genuine knowledge. The fact that treatment effects are small is prior information as real as the trial data. Ignoring it (via Jeffreys’ or uniform priors) is not objectivity—it is a specific, and poor, modeling choice. The Pólya-Gamma augmentation of Polson et al. [2013] removes the historical computational barrier to exact Bayesian inference for the logistic models most commonly used in trial analysis.
Sequential allocation follows because the posterior updates after every patient, and each update should inform the next decision. Thompson’s (1933) allocation rule is the simplest expression of this idea; the backward induction framework of Christen and Nakamura [2003] and Carlin et al. [1998] provides the optimal one, made tractable by the sufficient-statistic reduction. The predictive probability approximation of Marion et al. [2025] makes real-time monitoring computationally practical.
Decision-theoretic design follows because clinical trials involve multiple objectives—efficacy, toxicity, information gain, cost—that must be traded off explicitly. The utility-based framework of Thall and Cook [2004] demonstrates that this tradeoff can be formalized and optimized, producing designs that are simultaneously more ethical (fewer patients on inferior treatments) and more informative (better identification of optimal doses and subgroups).
The regulatory environment has evolved to accommodate these methods. The 2026 FDA guidance [Food and Drug Administration, 2026] provides a practical roadmap. The I-SPY 2 platform trial [Berry, 2025], the CALGB 49907 trial [Muss et al., 2009], and numerous adaptive designs in oncology and rare diseases demonstrate that Bayesian methods deliver on their theoretical promise in practice.

### v2

@@ 4.3 Adaptive enrichment
When a treatment benefits only a biomarker-defined subgroup, enrolling unselected patients wastes resources and exposes non-responders to risk. An enrichment design modifies enrolment criteria based on interim data, starting with broad enrolment and restricting to the responsive subgroup as evidence accumulates.
Formally, consider a trial with $G$ pre-specified subgroups $\mathcal{S}_{1},\ldots,\mathcal{S}_{G}$ defined by a biomarker. At each interim analysis after $n$ patients, the design computes subgroup-specific posterior probabilities $\Pr(\delta_{g}>0\mid D_{n})$ for each subgroup $g$, where $\delta_{g}$ is the treatment effect in subgroup $g$. The decision rule maps these probabilities to one of three actions: continue enrolling all subgroups, restrict enrolment to subgroups with $\Pr(\delta_{g}>0\mid D_{n})>\gamma_{\text{enrich}}$, or stop the trial entirely if no subgroup shows sufficient promise. The thresholds $\gamma_{\text{enrich}}$ and the stopping boundaries are calibrated by simulation to control the overall type I error rate while maximising power in the subgroup that benefits.
The decision at each interim—whether to continue broad enrolment, restrict, or stop—is a multi-armed bandit problem where the “arms” are enrolment strategies rather than treatments. This is precisely the structure addressed by the backward induction framework of Section 3: the state space is the collection of subgroup-specific sufficient statistics, and the utility function penalises both the cost of additional enrolment and the regret from treating non-responders.
The I-SPY 2 platform trial for neoadjuvant breast cancer provides a large-scale example. Patients are assigned to one of eight biomarker-defined subtypes, and each experimental arm is evaluated across up to ten molecular signatures (combinations of subtypes corresponding to possible clinical indications); arms that show $\geq 85\%$ Bayesian predictive probability of success in a confirmatory trial “graduate,” while arms that cross a futility boundary are dropped (Barker et al., 2009; Berry, 2025). This adaptive enrichment within a platform design has graduated nine of 23 experimental arms (39%) to phase III, with four of the nine graduates subsequently receiving marketing approval (Berry, 2025).
@@ 5.1 ECMO: the role of priors with minimal data
The 1985 trial of extracorporeal membrane oxygenation (ECMO) for newborns with persistent pulmonary hypertension (Bartlett et al., 1985) used a randomised play-the-winner design (Wei and Durham, 1978): adaptive randomisation that favours the more successful treatment. The first patient was randomised to ECMO and survived; the second to conventional medical therapy (CMT) and died; the remaining ten were all randomised to ECMO and all survived. Final tally: 11/11 ECMO survivors, 0/1 CMT survivors.
Ware (1989) discussed the statistical challenges; Kass and Greenhouse (1989) provided the Bayesian perspective. With a uniform prior $\text{Beta}(1,1)$ on each treatment’s survival probability, the posteriors are:
@@
The posterior probability $P(\theta_{\text{ECMO}}>\theta_{\text{CMT}}\mid\text{data})$ can be computed exactly via Thompson’s formula (8) and equals $90/91\approx 0.989$.
The critical issue is sensitivity to the prior. With only one control patient, the posterior for $\theta_{\text{CMT}}$ is dominated by the prior. Historical data suggested approximately 20% survival with CMT from large observational series. Encoding this as $\text{Beta}(4,16)$—with prior mean $4/20=0.20$ and effective sample size on the order of 20—yields $\theta_{\text{CMT}}\mid\text{data}\sim\text{Beta}(4,17)$ with posterior mean 0.19, and $P(\theta_{\text{ECMO}}>\theta_{\text{CMT}}\mid\text{data})$ remains above 0.99. The conclusion is robust because the ECMO evidence (11/11) is overwhelming; the prior matters only for the magnitude of the estimated difference, not the direction.
Berry (1989) emphasised that the Bayesian framework naturally handles the controversial stopping rule: under the likelihood principle, inference depends on the data actually observed, not on the stopping rule that generated them. A p-value for the same data depends on whether the investigator planned to stop after 12 patients, planned to stop after one CMT death, or had no fixed plan—the Bayesian posterior does not.
The binary backward induction of Section 3.5 provides a prospective design framework for trials like ECMO. Using a $\text{Beta}(4,16)$ prior for CMT (encoding the historical 20% survival rate with effective sample size 20) and a $\text{Beta}(1,1)$ prior for ECMO (noninformative for the experimental arm), we compute the calibrated backward induction policy (24) with $\gamma=0.975$, per-stage cost $c=0.001$, and horizon $T=100$. Figure 8 shows the resulting stopping boundaries projected onto the posterior mean treatment effect $\hat{\delta}_{k}$, together with simulated trial paths under the ECMO-like scenario.
Under the ECMO-like scenario ($p_{\text{ECMO}}=0.80$, $p_{\text{CMT}}=0.20$), the optimal design stops at a median of 2 patients per arm ($E[N]=3.0$), with 83% probability of declaring ECMO superior. Under a moderate effect ($p_{\text{ECMO}}=0.50$), the median rises to 6 patients per arm. Under the null ($p_{\text{ECMO}}=p_{\text{CMT}}=0.20$), the type I error is 0.13—elevated relative to the symmetric-prior simulations of Section 3.7 because the informative CMT prior concentrates control-arm uncertainty more quickly, increasing the chance that random fluctuations in the ECMO arm produce apparent superiority. This illustrates both the ethical advantage of decision-theoretic designs for large effects and the sensitivity of Bayesian sequential designs to prior asymmetry.
Figure 8: Optimal stopping regions from calibrated backward induction for the ECMO trial design, with $\text{Beta}(1,1)$ prior for ECMO and $\text{Beta}(4,16)$ prior for CMT. The asymmetric priors produce a wider stopping region for ECMO superiority (upper) than for CMT superiority (lower). Coloured curves show eight simulated trial paths under the ECMO-like scenario ($p_{\text{ECMO}}=0.80$, $p_{\text{CMT}}=0.20$); all terminate within 10 stages.
The question of ECMO efficacy was revisited in the EOLIA (Extracorporeal Membrane Oxygenation for Severe Acute Respiratory Distress Syndrome) trial, a separate randomised study of ECMO for adult ARDS (acute respiratory distress syndrome) (Goligher et al., 2018). The frequentist p-value was 0.09 (not significant at the conventional 0.05 level), but Bayesian analyses with various priors showed 88–99% posterior probability that ECMO reduces mortality—a substantially more informative summary than a binary significant/not-significant determination.
@@ 5.2 CALGB 49907: predictive probability monitoring
The Cancer and Leukemia Group B trial 49907 tested whether capecitabine could replace standard chemotherapy in women with early-stage breast cancer aged 65 or older (Muss et al., 2009). The planned sample size ranged from 600 to 1800 patients, with Bayesian predictive probability monitoring for futility and noninferiority at pre-specified interim analyses.
The noninferiority margin corresponded to a 5-year relapse-free survival of 53% for capecitabine versus 60% for standard therapy (equivalently, a hazard ratio of 1.24). At each interim analysis, the predictive probability was computed: given current data, what is the probability that future follow-up will yield a conclusive result?
At the first interim analysis after 600 patients, the hazard ratio (standard/capecitabine) was 0.53—equivalently, capecitabine patients had roughly $1/0.53\approx 1.9$ times the hazard of recurrence. The Bayesian posterior probability that capecitabine failed the noninferiority criterion exceeded 96%, surpassing the futility threshold of 80%. Enrolment was discontinued. After an additional year of follow-up, the hazard ratio (capecitabine/standard) was 2.09 (95% CI: 1.38–3.17; $P<0.001$), confirming that standard chemotherapy is superior.
The trial achieved its scientific goal with 633 patients instead of the 1800 maximum—a savings of 65% in sample size and years of accrual time. Berry (2025) describes CALGB 49907 as, to his knowledge, the first fully Bayesian randomised trial published in the New England Journal of Medicine. The design used noninformative priors for treatment effects; the Bayesian contribution was entirely in the decision framework (predictive probabilities for stopping), not in the priors.
@@ 5.3 Platform trials and the Bayesian Time Machine
Saville et al. (2022) proposed the Bayesian Time Machine: a hierarchical model that smooths control-arm response rates across time, allowing analyses to use all control data while adjusting for temporal drift. The model includes time effects as random effects with a smoothing prior that penalises abrupt changes, so that closer time intervals are modelled with similar response rates.
In simulations, the Bayesian Time Machine provides estimates with smaller mean squared error than either concurrent-only or pooled-controls analyses. The smoothing prior acts as a compromise: when temporal drift is absent, it borrows efficiently across all time periods; when drift is present, it limits borrowing to adjacent periods. The I-SPY 2 platform trial for neoadjuvant breast cancer (Barker et al., 2009) employed this approach across 23 experimental therapies, of which nine graduated to Phase III based on exceeding an 85% Bayesian predictive probability of confirmatory trial success (Berry, 2025).
@@ 6 The Regulatory Landscape
The FDA’s January 2026 draft guidance on Bayesian methodology (Food and Drug Administration, 2026) specifies concrete requirements for Bayesian submissions:
  1. (a)
    pre-specification of Bayesian success criteria and decision thresholds;
  2. (b)
    evaluation of operating characteristics through simulation, including type I error under the null hypothesis;
  3. (c)
    justification of prior distributions, with quantification of their influence (e.g., effective sample size);
  4. (d)
    sensitivity analyses examining robustness to prior specification and prior-data conflict;
  5. (e)
@@
The guidance identifies specific settings where Bayesian methods have supported successful submissions: borrowing from previous trials (e.g., REBYOTA for recurrent C. difficile), augmenting concurrent controls with external data (e.g., GBM AGILE for glioblastoma, Precision Promise for pancreatic cancer), paediatric extrapolation (e.g., empagliflozin, linagliptin for paediatric type 2 diabetes), and dose-finding trials in oncology.
The guidance can be read as accommodating two regulatory paths for Bayesian trials. The first requires demonstration of frequentist type I error control through simulation; the second may permit direct posterior probability thresholds provided that the prior and utility function are pre-specified and agreed upon with the agency, though to the authors’ knowledge no approved submission has yet relied solely on this path. Most successful Bayesian submissions to date have followed the first path, using predictive probability monitoring or prior borrowing calibrated to control type I error. The backward induction framework of this paper fits the second path: the utility function (11) and declaration threshold $\gamma$ are pre-specified, and the power frontier (Figure 7) provides exactly the simulation-based operating characteristics the guidance requires under items (a) and (b). Whether the inflated type I error inherent in utility-optimal designs (Section 3.7) is acceptable remains a regulatory judgement that will depend on the clinical context—rare diseases and paediatric extrapolation, where the guidance explicitly endorses Bayesian methods, are the most natural settings.
A common misconception equates Bayesian methods primarily with informative priors for borrowing. In fact, CALGB 49907 used noninformative priors (Muss et al., 2009), and I-SPY 2 used minimally informative treatment-effect priors overwhelmed by accumulating data (Berry, 2025). In both cases the Bayesian contribution was the decision framework—predictive probabilities for stopping and adaptive enrichment—not the priors. Predictive probabilities, while related to frequentist conditional power, answer a different question: given all observed data, what is the probability of eventual trial success?
@@ 7 Discussion
The central theme of this paper is that Bayesian methods can reach strong conclusions with small samples—and with informative priors, even smaller ones—though the sample-size savings come with trade-offs in frequentist operating characteristics that depend on the utility function and cost parameters (Section 3.7). The mechanism is the same in every example: the posterior precision equals the sum of prior precision and data precision, so each observation moves the posterior toward a decision boundary, and an informative prior gives a head start. When evidence accumulates in favour of one arm, the trial stops early; when evidence accumulates against it, the trial stops early for futility. Either way, patients are spared. The ECMO trial reached a definitive conclusion with 12 patients. CALGB 49907 stopped at 633 instead of 1800—a 65% saving. The multi-centre topical cream analysis produced 98.5% posterior probability of treatment efficacy from centres with as few as five patients per arm by borrowing strength across sites. In I-SPY 2, adaptive enrichment across biomarker-defined subtypes graduated nine of 23 experimental arms to Phase III, four of which have received marketing approval. In each case, the Bayesian framework converted prior knowledge and accumulating data into rapid decisions that a fixed-sample design could not have reached without enrolling far more patients.
These applied findings rest on three methodological pillars that are consequences of a single principle: probability is the language of uncertainty, and decisions under uncertainty should maximise expected utility. First, priors are not optional nuisance parameters but expressions of genuine knowledge. The fact that treatment effects are small is prior information as real as the trial data; ignoring it via Jeffreys’ or uniform priors is not objectivity but a specific modelling choice with known deficiencies (Section 2). The conditional means prior framework of Bedrick et al. (1996) provides a principled way to encode this knowledge for regression models, and the Pólya-Gamma augmentation of Polson et al. (2013) removes the historical computational barrier to exact Bayesian inference for the logistic models most commonly used in trial analysis. Second, the posterior updates after every patient, and each update should inform the next decision. Thompson’s (1933) allocation rule is the simplest expression of this idea; the backward induction framework of Christen and Nakamura (2003) and Carlin et al. (1998) provides the optimal rule under the assumed model and loss function, made tractable by the sufficient-statistic reduction.
The exact binary backward induction of Section 3.5 extends this to the two-arm setting most common in practice, while the Pólya-Gamma bridge (Section 3.6) connects it to covariate-adjusted logistic regression through a conditional Gaussian structure. Third, clinical trials involve multiple objectives—efficacy, toxicity, information gain, cost—that the utility-based framework of Thall and Cook (2004) trades off explicitly, producing designs that can improve both patient outcomes and statistical information.
The simulation study (Section 3.7) reveals a tension at the heart of Bayesian optimal design. The uncalibrated backward induction rule stops when the expected benefit of further sampling falls below the per-stage cost, achieving expected sample sizes of 14–26 patients per arm but with low power (0.03–0.18), because the small accumulated sample rarely pushes posterior probabilities past a stringent declaration threshold. The calibrated variant, which embeds the declaration threshold directly in the terminal utility, resolves much of this gap: at the same cost ($c=5\times 10^{-4}$), power at the largest treatment effect rises from 0.18 to 0.70 while expected sample size roughly doubles from 14 to 24 patients per arm—still well below the 48–100 required by competitors. Reducing the cost further to $c=10^{-4}$ pushes power to 0.81 at $E[N]=29$. However, both calibrated designs inflate type I error to 0.12–0.14 (versus 0.03 for the uncalibrated rule), because the sequential Bayesian policy provides more opportunities for the posterior probability to exceed $\gamma$ by chance under the null. The power frontier (Figure 7) makes the trade-off explicit and tunable: the per-stage cost $c$ controls the operating point on the E[N]-versus-power curve, and informative priors can partially mitigate type I error inflation (Table 3). In rare-disease or paediatric settings where each patient has high individual value, an operating point with low $E[N]$ may be preferred; in confirmatory trials seeking regulatory approval, predictive probability monitoring and group sequential designs offer tighter type I error control with power close to the fixed-sample benchmark. The principal contribution of the backward induction framework is therefore not that it outperforms these alternatives on frequentist criteria, but that it delivers the optimal policy for a stated utility—exact in the uncalibrated Beta-Binomial recursion, approximate in the calibrated variant where a Normal approximation to $\Pr(p_{1}>p_{0}\mid\text{data})$ gates the terminal reward—and makes the resulting trade-offs—between sample size, power, and error rate—explicit and tunable.
Backward induction is best understood as a design option on the power–sample-size frontier, suited to patient-sparing settings such as rare diseases, paediatrics, and early adaptive phases, rather than a replacement for confirmatory group sequential or fixed-sample designs. In regulatory practice, $c$ and $\gamma$ should be calibrated by simulation to meet pre-specified type I error and power targets before the trial begins, as required by the FDA guidance (Food and Drug Administration, 2026). The case studies of Section 5 carry the practical argument for Bayesian methods; the simulation reveals the theoretical limits of decision-theoretic optimality when judged by frequentist operating characteristics.
Several limitations should be noted. Our analysis draws on published trial summaries rather than patient-level data, so the case studies illustrate rather than independently replicate the original findings. The Normal model backward induction assumes known observation variance; unknown variance adds a second sufficient statistic and a correspondingly larger grid. The binary backward induction is exact on the integer lattice but scales as $O(T^{3})$, limiting practical horizons to a few hundred patients per arm; for larger trials, approximate methods such as predictive probability monitoring are more tractable. The Pólya-Gamma bridge to logistic regression (Section 3.6) relies on a Laplace approximation at the posterior mode of the augmented variables. Table 1 shows that this approximation is accurate across the sample sizes tested (mean absolute error below 0.01 for $n\geq 10$ per arm), though accuracy at very small samples ($n<10$) or with complete separation should be verified against exact calculations. The conjugate and logistic models that make Pólya-Gamma augmentation and conditional means priors tractable may not capture the complexities of time-to-event, longitudinal, or competing-risk endpoints that arise in modern platform trials. Finally, we do not address the practical challenges of prior elicitation from clinical teams—a step that remains the principal barrier to routine adoption of informative priors.
The 2026 FDA guidance (Food and Drug Administration, 2026) now provides a concrete regulatory pathway, and the accumulating record of Bayesian adaptive trials in oncology, rare diseases, and platform settings suggests that the remaining barriers are practical—prior elicitation, software infrastructure, and institutional familiarity—rather than methodological.
Data and code availability
No patient-level data were used in this paper. All figures are reproducible from the R scripts provided as supplementary material.
Competing interests
The authors declare no competing interests.
Author contributions
A.S. contributed clinical trial expertise and case study material. V.S. developed the mathematical framework, computational examples, and wrote the manuscript. N.P. provided the decision-theoretic perspective and editorial guidance.